PLEASE ANSWER THESE TWO QUESTIONS, I'M NEW AT THIS.
QUESTION 1- MEASURE THE DISTANCE OF THE DIAGONAL OF YOUR COMPUTER SCREEN TO THE NEAREST SIXTEENTH OF AN INCH, MEASURE THE HEIGHT OF THE SCREEN ALONG THE VERTICAL AS WELL. USE THE PYTHAGOREAN THEOREM TO FIND THE WIDTH ALONG THE HORIZOTAL.MY MEASUREMENTS ARE 21INCHES DIAGONAL AND 14 INCHES IN HEIGHT.
QUESTION 2- IF F(T)=2T^2-4T-1, FIND F(2) AND F(-1)

Answer by stanbon(75887)

QUESTION 1- MEASURE THE DISTANCE OF THE DIAGONAL OF YOUR COMPUTER SCREEN TO THE NEAREST SIXTEENTH OF AN INCH, MEASURE THE HEIGHT OF THE SCREEN ALONG THE VERTICAL AS WELL. USE THE PYTHAGOREAN THEOREM TO FIND THE WIDTH ALONG THE HORIZOTAL.MY MEASUREMENTS ARE 21INCHES DIAGONAL AND 14 INCHES IN HEIGHT.
-------------
21^2 = 14^2 + h^2
h^2 = 441-196
h = sqrt(245)
h = 15.6525 inches
----------------------
QUESTION 2- IF F(T)=2T^2-4T-1,
FIND F(2) = 2(2^2)-4*2-1 = 8-8+1 = 1
--------------------------
AND F(-1) = 2(-1)^2-4*-1-1 = 2+4-1 = 5
================
Cheers,
Stan H.